# AutoAgents: A Framework for Automatic Agent Generation

### v1

@@ Abstract
Large language models (LLMs) have enabled remarkable advances in automated task-solving with multi-agent systems. However, most existing LLM-based multi-agent approaches rely on predefined agents to handle simple tasks, limiting the adaptability of multi-agent collaboration to different scenarios. Therefore, we introduce AutoAgents, an innovative framework that adaptively generates and coordinates multiple specialized agents to build an AI team according to different tasks. Specifically, AutoAgents couples the relationship between tasks and roles by dynamically generating multiple required agents based on task content and planning solutions for the current task based on the generated expert agents. Multiple specialized agents collaborate with each other to efficiently accomplish tasks. Concurrently, an observer role is incorporated into the framework to reflect on the designated plans and agents’ responses and improve upon them. Our experiments on various benchmarks demonstrate that AutoAgents generates more coherent and accurate solutions than the existing multi-agent methods. This underscores the significance of assigning different roles to different tasks and of team cooperation, offering new perspectives for tackling complex tasks. The repository of this project is available at https://github.com/LinkSoul-AI/AutoAgents.
@@ 1 Introduction
Large language models (LLMs) have exhibited astounding capabilities as versatile task-solving agents, endowed with a rich blend of knowledge and skills. Nevertheless, they still face difficulties (Qin et al., 2023; OpenAI, 2023; Bubeck et al., 2023) in tackling various tasks that require intensive knowledge and reasoning, such as avoiding hallucination (Maynez et al., 2020), employing slow-thinking strategies (Sloman, 1996), ensuring trustworthiness (Wang et al., 2023a), and in combining diverse domain knowledge and long-horizon planning. In contrast, humans often exploit the benefits of collaborative problem solving, which enables them to work together effectively to solve non-routine problems in diverse domains and enhance the quality and reliability of the solutions by distributing the workload among specialties and applying a diversity of perspectives and expertise (Nelson, 2013; Roschelle & Teasley, 1995; Barron, 2000).
Inspired by collaborative problem solving, several recent works (Wang et al., 2023c; Du et al., 2023; Liang et al., 2023; Hao et al., 2023) have improved the task-solving capabilities of LLMs by integrating multi-agent discussion. However, most of these multi-agent systems depend on handcrafted or user-specified agents, with specific roles and necessitating human supervision, which often restricts the scope of collaborative applications. Moreover, manually creating a large number of experts often consumes a lot of resources. In order to adaptively solve more complex problems, this paper aims to explore a method of adaptively generating task experts and completing different tasks through multi-level collaborative cooperation among multiple experts.
@@
Framework Generate Agent Dynamically Number of Agent Multi-agent Conversation Self-Refinement Agent Collaborative Refinement Action AutoGPT (Gravitas, 2023) 1 BabyAGI (Nakajima, 2023) 3 Generative Agents (Park et al., 2023) 25 Camel (Li et al., 2023) 2 GPT-bargaining (Fu et al., 2023) 3 MetaGPT (Hong et al., 2023) Unlimited AutoGen (Wu et al., 2023) Unlimited Social Simulacra (Park et al., 2022) Unlimited Epidemic Modeling (Williams et al., 2023) Unlimited ExpertPrompting (Xu et al., 2023) 1 SSP (Wang et al., 2023c) Unlimited AgentVerse (Chen et al., 2023a) Unlimited AutoAgents Unlimited
@@ 2 Related Work
LLM-based Autonomous Agents. LLMs have been widely used as core controllers for autonomous agents that can accomplish specific objectives. Auto-GPT (Gravitas, 2023) is an early work that leverages an LLM as an AI agent that can autonomously achieve a given goal with the help of several tools. However, Auto-GPT does not support multi-agent collaboration and can only work in isolation. One way to enhance the task-solving capabilities of LLMs is to assign different roles and responsibilities to multiple LLMs and let them coordinate their actions to achieve a common goal. For example, BabyAGI (Nakajima, 2023) is an AI-powered task management system with multiple LLM-based agents. One agent creates new tasks based on the previous task’s objective and result, another agent prioritizes the task list, and another agent completes tasks. BabyAGI is a multi-agent system with a fixed order of agent communication. MetaGPT (Hong et al., 2023) is a multi-agent framework for assigning different roles to GPTs to form a collaborative software entity for complex tasks. It is a specialized LLM-based multi-agent framework for collaborative software development. Camel (Li et al., 2023) is an LLM-based communicative agent framework that demonstrates how role-playing can be used to enable chat agents to communicate with each other for task completion. However, Camel does not support tool-using. Several recent works (Wang et al., 2023c; Du et al., 2023; Liang et al., 2023; Hao et al., 2023) have enhanced the task-solving capabilities of LLMs by integrating multi-agent discussion. For instance,  Wang et al. (2023c) proposes a multi-agent debate system that allows LLMs to argue for or against a given claim and generate a debate summary.  Du et al. (2023) introduce a multi-agent dialogue system that enables LLMs to exchange information and opinions on a given topic and generate a dialogue report. AutoGen (Wu et al., 2023) is a framework that enables the development of LLM applications using multiple agents that can converse with each other to solve tasks. However, most of these multi-agent systems rely on handcrafted or user-specified agents with specific roles and do not support the automatic generation of agents, which often limits the scope of collaborative applications.
@@
Agent Generalization. Several studies  (Park et al., 2022; Williams et al., 2023) employ LLMs to generate agents for social simulacra and epidemic modeling, demonstrating how this technique can facilitate designers in assessing and improving their modeling designs prior to deploying them to real users. Likewise, ExpertPrompting (Xu et al., 2023) devised a method to generate diverse profiles of agents that can cooperate with human users to accomplish tasks with minimal supervision. However, this method still depends on a restricted set of predefined agents, and the generated agents vary only in their profiles, not in their problem-solving capacities. Recently, SSP (Wang et al., 2023c) and AgentVerse (Chen et al., 2023a) have proposed frameworks for automatically generating unlimited agents. SSP enables LLMs to generate agents for problem input by providing some agent samples, and has these agents solve the problem. AgentVerse generates the execution plan through the generated agents’ discussions and adds evaluation strategies for cyclic execution. Unlike the previous two methods, AutoAgents pays more attention to the details of the generated agents and plans, and improves the efficiency of task execution by employing collaborative refinement actions and self-refinement agents, as illustrated in Table 1.
@@ 3.1 Drafting Stage
Empirical evidence  (Woolley et al., 2015) suggests that diversity within human groups fosters diverse perspectives, which enhances the group’s performance across various tasks. The drafting stage, which determines the composition of a multi-agent group, plays a crucial role in setting the upper limits of the group’s capabilities. Therefore, it is imperative to generate the optimal agent team and execution plan that can maximize the group’s potential.
Predominant methodologies (Gravitas, 2023; Hong et al., 2023; Wu et al., 2023) for assigning role descriptions to autonomous agents rely heavily on human intuition and prior knowledge, requiring manual assignment based on task understanding. Consistent with several parallel findings (Xu et al., 2023; Wang et al., 2023c; Chen et al., 2023a), dynamically designing agents with different roles can significantly enhance their efficacy. However, the scalability and rationality of agent and plan generation are still unclear, especially in the face of various complex problem environments.
@@
Plan Generation. In parallel to agent generation, the Planner formulates the execution plan and promotes its progressive improvement through reciprocal communication with the Plan Observer. For a given task, the Planner delineates the specific steps $\{\mathcal{S}_{1},\mathcal{S}_{2},\cdots\mathcal{S}_{n}\}$ to accomplish it in the execution plan $\mathcal{P}$. Each step $\mathcal{S}_{i}$ entails a clear identification of the agent $\mathcal{A}_{j}$ responsible for it, as well as the input information and expected output required for it.
The Plan Observer subsequently validates the execution plan $\mathcal{P}=\{\mathcal{S}_{1},\mathcal{S}_{2},\cdots\mathcal{S}_{n}\}$ according to the agent list $\{\mathcal{A}_{1},\mathcal{A}_{2},\cdots,\mathcal{A}_{n}\}$ and the task content. It firstly ensures that each step has a corresponding agent and that the step content is coherent and concise. It secondly assesses whether all the steps are sufficient, whether the task can be accomplished, and whether there are any gaps that need to be filled. It finally provides feedback to the Planner, who further refines the execution plan accordingly. After $n$ rounds of dialogue between the Planner and the Plan Observer, the ultimate execution plan for achieving the task is established.
@@ 3.2 Execution Stage
In the drafting phase, the framework generates an agent list and an execution plan based on the task requirements. Then, the framework creates corresponding roles and executes the plans in the execution environment [^1] Execution environment of AutoAgents is built based on MetaGPT’s environment and workspace (Hong et al., 2023).. The communication and cooperation among multi-agent systems are essential for accomplishing the tasks effectively. This section elaborates on the communication among multiple agents, the task execution strategies, and the knowledge-sharing mechanisms.
Communication of Multiple Agent. The communication structures among agents have been investigated by many studies (Chen et al., 2023a; Wang et al., 2023c; Qian et al., 2023; Chan et al., 2023) to examine their impact on task performance. In this framework, we adopt the vertical communication paradigm, which assigns different responsibilities to agents according to their roles. To facilitate the specific division of labor among the agents in the generated team, we introduce a predefined Action Observer as the team leader to coordinate the execution plan. Specifically,
@@
Self-refinement Agent. Besides the inter-agent communication, the performance of a single agent also exerts a significant impact on the overall quality of feedback results. Hence, drawing on mechanisms such as AutoGPT (Gravitas, 2023) and ReAct (Yao et al., 2022b), we have devised a self-refinement mechanism for an individual agent.
For a single agent $\mathcal{A}$, the action at step $t$ is at $a_{t}=l_{t}\cup p_{t}\cup o_{t}$, where $l_{t}$, where $l_{t}$ denotes the thought or the reasoning trace in the language space, which does not alter the external environment, and thus yields no observational feedback, $p_{t}$ represents the execution plan for better task completion, $o_{t}$ comprises the completion steps and execution output for this time.
@@
Knowledge Sharing Mechanism. AutoAgents also facilitates the sharing of execution results among various agents for improved communication and feedback. However, when the number of agents is large and a single agent has more self-iterations, it will generate more historical information. Due to the token limitation of LLM models, they often cannot encompass all information. Hence, this framework provides short-term memory, long-term memory, and dynamic memory based on agent units.
Short-term memory primarily focuses on a single agent and encompasses intermediate ideas, plans, and execution results generated during the self-refinement process of an individual agent. It is noteworthy that the self-refinement of a single agent often culminates in a summary of vital information in the final step of the self-refinement process.
Long-term memory mainly targets the communication between multiple agents, chiefly recording the tasks executed by an individual agent and the summary of crucial feedback information. This is essential for assessing the overall degree of task completion.
Dynamic memory mainly caters to agents with special needs. The Action Observer can access all short-term and long-term memories, dynamically extracting supplementary information from short-term and long-term memories based on the information required by the agent to execute tasks, aiding in enhancing the task execution efficiency of a single agent.
@@ 4.1 Open-ended Question Answer
Task Description. Open-ended Question Answering is a crucial and challenging task in the domain of NLP and generative AI. It requires an AI system to produce coherent, elaborate, and human-like responses to questions that have no predetermined or fixed set of possible answers.  (Zheng et al., 2023) proposed MT-bench, a benchmark consisting of 80 high-quality collected open-ended questions from various categories such as common sense, counterfactual, coding, etc. We then utilize AutoAgents to produce collaborative answers based on multiple generated agents and compare them with the responses given by Vicuna-13B, ChatGPT, and GPT-4.
Table 2: Win Rate of AutoAgents over other models on Open-ended Question Answer, with FairEval (Wang et al., 2023b) and HumanEval serving as evaluators.
Evaluator v.s. ChatGPT v.s. Vicuna-13B v.s. GPT-4 FairEval (Wang et al., 2023b) 96.3% 96.3% 76.3% HumanEval 75% 75% 62.5%
Evaluation Metrics. To measure the quality of open-ended responses with minimal evaluation bias, we adopt FairEval (Wang et al., 2023b) and HumanEval as the evaluation metrics for both the single agent and AutoAgents. FairEval incorporates several methods to mitigate the impact of various sources of bias, resulting in a better alignment with human judgment. For HumanEval, we enlist several volunteers to rate the responses from different models based on their helpfulness, reliability, accuracy, and level of detail.
Results. Table 2 demonstrates that AutoAgents outperforms individual LLM models in both FairEval based on LLM and Human evaluations. AutoAgent can produce more comprehensive and nuanced answers to open questions by synthesizing multiple expert models. It can also provide more elaborate explanations and justifications for its answers. More examples are given in the Appendix 4.4.

### v3

@@ Abstract
Large language models (LLMs) have enabled remarkable advances in automated task-solving with multi-agent systems. However, most existing LLM-based multi-agent approaches rely on predefined agents to handle simple tasks, limiting the adaptability of multi-agent collaboration to different scenarios. Therefore, we introduce AutoAgents, an innovative framework that adaptively generates and coordinates multiple specialized agents to build an AI team according to different tasks. Specifically, AutoAgents couples the relationship between tasks and roles by dynamically generating multiple required agents based on task content and planning solutions for the current task based on the generated expert agents. Multiple specialized agents collaborate with each other to efficiently accomplish tasks. Concurrently, an observer role is incorporated into the framework to reflect on the designated plans and agents’ responses and improve upon them. Our experiments on various benchmarks demonstrate that AutoAgents generates more coherent and accurate solutions than the existing multi-agent methods. This underscores the significance of assigning different roles to different tasks and of team cooperation, offering new perspectives for tackling complex tasks. The repository of this project is available at https://github.com/Link-AGI/AutoAgents.
@@ 1 Introduction
Large language models (LLMs) have exhibited astounding capabilities [26, 22, 29] as versatile task-solving agents, endowed with a rich blend of knowledge and skills. Nevertheless, they still face difficulties [26, 22, 2] in tackling various tasks that require intensive knowledge and reasoning, such as avoiding hallucination [19], employing slow-thinking strategies [30], ensuring trustworthiness [32], and in combining diverse domain knowledge and long-horizon planning. In contrast, humans often exploit the benefits of collaborative problem solving, which enables them to work together effectively to solve non-routine problems in diverse domains and enhance the quality and reliability of the solutions by distributing the workload among specialties and applying a diversity of perspectives and expertise [21, 27, 1].
Inspired by collaborative problem solving, several recent works [34, 7, 17, 11] have improved the task-solving capabilities of LLMs by integrating multi-agent discussion. However, most of these multi-agent systems depend on handcrafted or user-specified agents, with specific roles and necessitating human supervision, which often restricts the scope of collaborative applications. Moreover, manually creating a large number of experts often consumes a lot of resources. In order to adaptively solve more complex problems, this paper aims to explore a method of adaptively generating task experts and completing different tasks through multi-level collaborative cooperation among multiple experts.
@@
Framework Dynamic Agent Generation Method Number of Agent Multi-agent Conversation Self-Refinement Action Collaborative Refinement Action AutoGPT [10] 1 BabyAGI [20] 3 Generative Agents [23] 25 Camel [16] 2 GPT-bargaining [8] 3 MetaGPT [12] Unlimited AutoGen [37] Unlimited Social Simulacra [24] Single Agent Unlimited Epidemic Modeling [35] Single Agent Unlimited ExpertPrompting [38] Single Agent 1 SSP [34] Single Agent Unlimited AgentVerse [5] Single Agent Unlimited AutoAgents Multi-agent Discussion Unlimited
@@ 2 Related Work
LLM-based Autonomous Agents. LLMs have been widely used as core controllers for autonomous agents that can accomplish specific objectives. Auto-GPT [10] is an early work that leverages an LLM as an AI agent that can autonomously achieve a given goal with the help of several tools. However, Auto-GPT does not support multi-agent collaboration and can only work in isolation. One way to enhance the task-solving capabilities of LLMs is to assign different roles and responsibilities to multiple LLMs and let them coordinate their actions to achieve a common goal. For example, BabyAGI [20] is an AI-powered task management system with multiple LLM-based agents. One agent creates new tasks based on the previous task’s objective and result, another agent prioritizes the task list, and another agent completes tasks. BabyAGI is a multi-agent system with a fixed order of agent communication. MetaGPT [12] is a multi-agent framework for assigning different roles to GPTs to form a collaborative software entity for complex tasks. It is a specialized LLM-based multi-agent framework for collaborative software development. Camel [16] is an LLM-based communicative agent framework that demonstrates how role-playing can be used to enable chat agents to communicate with each other for task completion. However, Camel does not support tool-using. Several recent works [34, 7, 17, 11, 31, 15] have enhanced the task-solving capabilities of LLMs by integrating multi-agent discussion. For instance,  [34] proposes a multi-agent debate system that allows LLMs to argue for or against a given claim and generate a debate summary.  [7] introduce a multi-agent dialogue system that enables LLMs to exchange information and opinions on a given topic and generate a dialogue report. AutoGen [37] is a framework that enables the development of LLM applications using multiple agents that can converse with each other to solve tasks. However, most of these multi-agent systems rely on handcrafted or user-specified agents with specific roles and do not support the automatic generation of agents, which often limits the scope of collaborative applications.
@@
Agent Generalization. Several studies  [24, 35] employ LLMs to generate agents for social simulacra and epidemic modeling, demonstrating how this technique can facilitate designers in assessing and improving their modeling designs prior to deploying them to real users. Likewise, ExpertPrompting [38] devised a method to generate diverse profiles of agents that can cooperate with human users to accomplish tasks with minimal supervision. However, this method still depends on a restricted set of predefined agents, and the generated agents vary only in their profiles. Recently, SSP [34] and AgentVerse [5] have proposed frameworks for automatically generating unlimited agents. SSP enables LLMs to generate agents for problem input by providing some agent samples, and has these agents solve the problem. AgentVerse generates the execution plan through the generated agents’ discussions and adds evaluation strategies for cyclic execution. Unlike the previous two methods, AutoAgents places a heightened emphasis on the reliability of its generated agents and strategic plans, thereby enhancing task execution effect through the utilization of collaborative refinement actions and the integration of self-refinement actions , as illustrated in Table 1.
@@ 3.1 Drafting Stage
Empirical evidence  [36] suggests that diversity within human groups fosters diverse perspectives, which enhances the group’s performance across various tasks. The drafting stage, which determines the composition of a multi-agent group, plays a crucial role in setting the upper limits of the group’s capabilities. Therefore, it is imperative to generate the optimal agent team and execution plan that can maximize the group’s potential.
Predominant methodologies [10, 12, 37] for assigning role descriptions to autonomous agents rely heavily on human intuition and prior knowledge, requiring manual assignment based on task understanding. Consistent with several parallel findings [38, 34, 5], dynamically designing agents with different roles can significantly enhance their efficacy. However, the scalability and rationality of agent and plan generation are still unclear, especially in the face of various complex problem environments.
@@
Plan Generation. In parallel to agent generation, the Planner formulates the execution plan and promotes its progressive improvement through reciprocal communication with the Plan Observer. For a given task, the Planner delineates the specific steps $\{\mathcal{S}_{1},\mathcal{S}_{2},\cdots\mathcal{S}_{n}\}$ to accomplish it in the execution plan $P$. Each step $\mathcal{S}_{i}$ entails a clear identification of the agent $\mathcal{A}_{j}$ responsible for it, as well as the input information and expected output required for it.
The Plan Observer subsequently validates the execution plan $P=\{\mathcal{S}_{1},\mathcal{S}_{2},\cdots\mathcal{S}_{n}\}$ according to the agent list $\{\mathcal{A}_{1},\mathcal{A}_{2},\cdots,\mathcal{A}_{n}\}$ and the task content. It firstly ensures that each step has a corresponding agent and that the step content is coherent and concise. It secondly assesses whether all the steps are sufficient, whether the task can be accomplished, and whether there are any gaps that need to be filled. It finally provides feedback to the Planner, who further refines the execution plan accordingly. After $n$ rounds of dialogue between the Planner and the Plan Observer, the ultimate execution plan for achieving the task is established.
@@ 3.2 Execution Stage
In the drafting phase, the framework generates an agent list and an execution plan based on the task requirements. Then, the framework creates corresponding roles and executes the plans in the execution environment [^1] Execution environment of AutoAgents is built based on MetaGPT’s environment and workspace [12].. The communication and cooperation among multi-agent systems are essential for accomplishing the tasks effectively. This section elaborates on the communication among multiple agents, the task execution strategies, and the knowledge-sharing mechanisms.
Communication of Multiple Agent. The communication structures among agents have been investigated by many studies [5, 34, 25, 3] to examine their impact on task performance. In this framework, we adopt the vertical communication paradigm, which assigns different tasks to agents according to their roles. To facilitate the specific division of labor among the agents in the generated team, we introduce a predefined Action Observer as the team leader to coordinate the execution plan. Specifically,
@@
Self-refinement Agent. Besides the inter-agent communication, the performance of a single agent also exerts a significant impact on the overall quality of feedback results. Hence, drawing on mechanisms such as AutoGPT [10] and ReAct [40], we have devised a self-refinement mechanism for an individual agent.
For a single agent $\mathcal{A}$, the action at step $t$ is at $a_{t}=l_{t}\cup p_{t}\cup o_{t}$, where $l_{t}$ denotes the thought or the reasoning trace in the language space, which does not alter the external environment, and thus yields no observational feedback, $p_{t}$ represents the execution plan for task completion, $o_{t}$ comprises the completion steps and execution output for this time.
@@
Figure 4: Legend of Three Knowledge Sharing Mechanisms. (a) Long-term memory focuses on chronicling the historical trajectory of multiple actions. (b) Short-term memory records the history of the self-refinement or collaborative refinement phases of an individual action. (c) Dynamic memory serves actions necessitating specialized attention extracted from the long-term memory.
Knowledge Sharing Mechanism. AutoAgents also facilitates the sharing of execution results among various agents for improved communication and feedback. However, when the number of agents is large and a single agent has more self-iterations, it will generate more historical information. Due to the token limitation of LLM models, they often cannot encompass all information. Hence, this framework provides short-term memory, long-term memory, and dynamic memory.
Short-term memory is chiefly concentrated on a singular action, encompassing the gamut of intermediary notions, strategies, and outcomes that emerge during the self-refinement or collaborative refinement phases of an individual action. It is salient to note that these actions frequently culminate in a distilled summary of critical information, epitomizing the final phase of the refinement trajectory.
Long-term memory principally focuses on chronicling the historical trajectory of multifarious actions, predominantly documenting the executed results of each task along with the synthesis of vital feedback information. This aspect is imperative for evaluating the comprehensive extent of task completion.
Dynamic memory predominantly serves actions necessitating specialized attention. The Action Observer, having access to long-term memory archives, adeptly extracts ancillary information, dynamically tailoring it to the specific requirements of the action for task execution. This process significantly augments the efficiency of a single action in task fulfillment.
0:  User task/Question
0:  Task solution/Answer
1:  Drafting Stage
2:  Initialize Planner $\mathcal{P}$, Agent Observer $\mathcal{O}_{\text{agent}}$, and Plan Observer $\mathcal{O}_{\text{plan}}$.
3:  $\mathcal{P}$ generates initial agent team $\{\mathcal{A}_{1},\mathcal{A}_{2},\cdots,\mathcal{A}_{n}\}$ and execution plan $P=\{\mathcal{S}_{1},\mathcal{S}_{2},\cdots\mathcal{S}_{n}\}$.
4:  repeat
5:     $\mathcal{O}_{\text{agent}}$ provides feedback on agent team.
6:     $\mathcal{P}$ refines agent team based on feedback.
7:     $\mathcal{O}_{\text{plan}}$ provides feedback on execution plan.
8:     $P$ refines execution plan based on feedback.
9:  until No feedback or reached the maximum iteration limit.
10:  Execution Stage:
11:  Initialize Action Observer $\mathcal{O}_{\text{action}}$ and long-term memory $M_{L}$.
12:  for $\{\mathcal{S}_{1},\mathcal{S}_{2},\cdots\mathcal{S}_{n}\}$ do
13:     $\mathcal{O}_{\text{action}}$ generate dynamic memory $M_{D}$.
14:     $\mathcal{O}_{\text{action}}$ assign task $\mathcal{S}_{k}$ and $M_{D}$ to corresponding agents $\{\mathcal{A}_{i},\cdots,\mathcal{A}_{j}\}$.
15:     Initialize short-term memory $M_{S}$.
16:     repeat
17:        for $\{\mathcal{A}_{i},\cdots,\mathcal{A}_{j}\}$ do
18:           Agent $\mathcal{A}_{m}$ analyzes $\mathcal{S}_{k}$, $M_{S}$ and $M_{D}$.
19:           Agent $\mathcal{A}_{m}$ plans the current step and executes this step.
20:           The execution result is stored in $M_{S}$.
21:        end for
22:     until No step or reached the maximum iteration limit.
23:     The execution results of task $\mathcal{S}_{k}$ are stored in $M_{L}$.
24:     $\mathcal{O}_{\text{action}}$ coordinates $\{\mathcal{S}_{1},\mathcal{S}_{2},\cdots\mathcal{S}_{n}\}$ and monitors execution.
25:  end for
26:  return Execution results of final step.
Algorithm 1 AutoAgents Execution Process.
@@ 4.1 Open-ended Question Answer
Task Description. Open-ended Question Answering is a crucial and challenging task in the domain of NLP and generative AI. It requires an AI system to produce coherent, elaborate, and human-like responses to questions that have no predetermined or fixed set of possible answers.  [41] proposed MT-bench, a benchmark consisting of 80 high-quality collected open-ended questions from various categories such as common sense, counterfactual, coding, etc. We then utilize AutoAgents to produce collaborative answers based on multiple generated agents and compare them with the responses given by Vicuna-13B, ChatGPT, and GPT-4.
Table 2: Win Rate of AutoAgents over other models on Open-ended Question Answer, with FairEval [33] and HumanEval serving as evaluators.
Evaluator v.s. ChatGPT v.s. Vicuna-13B v.s. GPT-4 FairEval [33] 96.3% 96.3% 76.3% HumanEval 75% 75% 62.5%
Evaluation Metrics. To measure the quality of open-ended responses with minimal evaluation bias, we adopt FairEval [33] and HumanEval as the evaluation metrics for both the single agent and AutoAgents. FairEval incorporates several methods to mitigate the impact of various sources of bias, resulting in a better alignment with human judgment. For HumanEval, we enlist several volunteers to rate the responses from different models based on their helpfulness, reliability, accuracy, and level of detail.
Results. Table 2 demonstrates that AutoAgents outperforms individual LLM models in both FairEval based on LLM and Human evaluations. AutoAgent can produce more comprehensive and nuanced answers to open questions by synthesizing multiple expert models. It can also provide more elaborate explanations and justifications for its answers. More examples are given in the Appendix A.